A data set includes 105 body temperatures of healthy adult humans for which (line over top of x)x=98.7 degrees F and s=0.64 degrees F.
Complete parts (a) and (b) below. SHOW WORK
a) What is the best point estimate of the mean body temperature of all healthy humans?
The best point estimate is______ degrees F(Type an interger or a decmial)
b)Using the sample statistics, construct a 99% confidence interval estimate of the mean body temperature of all healthy humans. Do the confidence interval limits contains 98.6 degrees F? What does the sample sugesst about the use of 98.6 degrees F as the mean body temperature?
What is the confidence interval estimate of the population mean u?
_______ Do the confidence interval limits contain 98.6 degrees F? No or Yes
What does this suggest about the use of 98.6 degrees F as the mean body Temperature?
A) This suggests that the mean body temperature could very possibly be 98.6 degrees F.
B) This suggest that the mean body temperature could be lower then 98.6 degrees F
C)This suggest that the mean body temperature could be higher then 98.6 degrees F
PLEASE HELP ME!!! I AS SO LOST AND DO NOT KNOW WHERE TO BEGIN PLEASE.

Found 2 solutions by stanbon, rothauserc:
Answer by stanbon(75887)

A data set includes 105 body temperatures of healthy adult humans for which (line over top of x)x=98.7 degrees F and s=0.64 degrees F.
Complete parts (a) and (b) below. SHOW WORK
a) What is the best point estimate of the mean body temperature of all healthy humans?
The best point estimate is x-bar = 98.7 degrees F(Type an interger or a decmial)
-----
b)Using the sample statistics, construct a 99% confidence interval estimate of the mean body temperature of all healthy humans.
sample mean = 98.7
Margin of Error = 2.5758 = 0.1609
99% CI:: 98.7-0.1609 < u < 98.7+0.1609
98.54 < u < 98.86
---------------------------------------------
Do the confidence interval limits contains 98.6 degrees F? Yes
----------------------
What does the sample suggest about the use of 98.6 degrees F as the mean body temperature?
We can be 99% confident that the mean body temp is 98.6
----------------------
What is the confidence interval estimate of the population mean u?
-----
_______ Do the confidence interval limits contain 98.6 degrees F? No or Yes
What does this suggest about the use of 98.6 degrees F as the mean body Temperature?
A) This suggests that the mean body temperature could very possibly be 98.6 degrees F.
B) This suggest that the mean body temperature could be lower then 98.6 degrees F
C)This suggest that the mean body temperature could be higher then 98.6 degrees F
--------------
Cheers,
Stan H.
------------

Answer by rothauserc(4718)

a) The best point estimate of the mean body temperature of all healthy humans is the mean (x with bar above it) of the given sample. The answer is 98.7 F
Note the sample mean provides the best estimate for the population mean
*******************************************************************************
b) steps to find 99% confidence interval
1. alpha = 1 - (confidence level / 100) = 1 - .99 = 0.01
2. critical probability(p*) P* = 1 - alpha/2 = 1 - 0.01/2 = 0.995
since the sample population size is > 30, we can use the z-tables to find the z-score associated with p*
z-score is 2.57
3. standard error of the mean is sample std dev / square root(sample size)
standard error of the mean = 0.64 / square root(105) = 0.062457605 = 0.06
4. Margin of error(ME) = critical value * standard error
ME = 2.57 * 0.06 = 0.1542 approx 0.15
5. confidence interval 98.7 + or - 0.15 (98.55, 98.85)
6. Yes, confidence interval includes 98.6 F
7. Suppose we took more samples of 105 people, since 98.6 F is include in our confidence interval, we can expect that 99% of such samples will include an average of 98.6 F.
8. 99% confidence interval of the population mean is (98.55, 98.85)
Therefore the answer is A.
*****************************************************************************